OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)?

Mehek (mehek14):

do you know the slope formula?

OpenStudy (anonymous):

yes, but the options for my answer are really confusing me they are: y = 4x − 2 y = 2x − 4 y = −3(x − 4) y = 3(x + 4)

OpenStudy (anonymous):

i keep getting y = -3x + 12

Mehek (mehek14):

you want to know why?

Mehek (mehek14):

do you know how to use distributive property?

Mehek (mehek14):

@cocomonga22 ?

OpenStudy (anonymous):

yes i want to know

OpenStudy (anonymous):

oh i get it now

Mehek (mehek14):

do you know the distributive property?

OpenStudy (anonymous):

wow i feel so stupid, its y= -3 (x-4)

Mehek (mehek14):

haha yes that's correct

OpenStudy (anonymous):

thanks! can i ask one more question please?

Mehek (mehek14):

sure ^_^

OpenStudy (anonymous):

how do you solve this problem? Solve the following for y: −6x + 2y = 8

Mehek (mehek14):

you have options?

OpenStudy (anonymous):

yes- y = −6x + 8 y = 6x + 8 y = −3x + 4 y = 3x + 4

Mehek (mehek14):

\(-6x+2y=8\) first you want to move -6x to move it, you add +6x that gives you \(2y=8+6x\) next you want to isolate y and get rid of 2 to do that, divide 2 by each number 2/2 = 0 8/2=4 6x/2=3x \(y=4+3x\)

OpenStudy (anonymous):

oh ok that makes sense now . Thanks for all your help! much appreciated

Mehek (mehek14):

anytime ^_^

OpenStudy (anonymous):

can i ask you one more question then?

OpenStudy (anonymous):

nevermind, thans for your help again!